According to a study done last year, the average monthly expenses for mobile phone loads of senior high school students in Sta. Rosa was P350.00. A statistics student believes that this amount has increased since January of this year. Is there a reason to believe that this amount has really increased if a random sample of 60 students has an average monthly expenses for mobile phone loads of P380.00? Use a 0.05 level of significance. Assume that the population standard deviation is P77.00.


1
Expert's answer
2021-05-07T10:37:51-0400

Let, Null hypothesis "H_0" : "\mu>P350"

Alternative hypothesis "H_{\alpha}" : "\mu\leq350"


Test statistic:


"z=\dfrac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{380-350}{77/\sqrt{60}}=3.02"


P-value:

"p=P(z>3.02)=1-0.9987=0.0013"


Since the P-value is less than 0.05, reject the null hypothesis.

So, we can conclude that monthly expenses for cell phone loads was not really increased since January of this year.